# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 7.2 Confirming points for the final states of apparatuses
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{j}\in\Omega_{j}$, every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exists a vector $|\Theta_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
@@
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the states of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
@@
The notion of the confirming point for the states of the system $\mathcal{S}$ is defined as follows: Let $V$ be the set of all $j\in\{1,\dots,n\}$ such that the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V\neq\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the system $\mathcal{S}$ itself if $j=1$ and the apparatus $\mathcal{A}_{j-1}$ otherwise, where $j:=\min V$. If $V=\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the apparatus $\mathcal{A}_{n}$. ∎
@@
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{i}\in\Omega_{i}$, every $m_{j}\in\Omega_{j}$, every $|\Psi_{L}\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exist a set $\{|\Theta_{L}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Theta_{R}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
@@
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
@@
Let $i\in\{1,\dots,n\}$. The notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as follows: Let $V_{i}$ be the set of all $j\in\{i+1,\dots,n\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V_{i}\neq\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{j-1}$ with $j:=\min V_{i}$. If $V_{i}=\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{n}$. Thus, the confirming point for the final states of the apparatus $\mathcal{A}_{n}$ is the apparatus $\mathcal{A}_{n}$ itself. ∎
Consider the case where there are only two apparatuses $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with $n=2$ and the final states of the apparatus $\mathcal{A}_{1}$ is not unchanged after the measurement by the apparatus $\mathcal{A}_{2}$. In this case, we cannot think the measurement $\mathcal{M}^{1}$ by the apparatus $\mathcal{A}_{1}$ to be completed immediately after the measurement $\mathcal{M}^{1}$ itself, since the state of the apparatus $\mathcal{A}_{1}$ is disturbed by the subsequent measurement $\mathcal{M}^{2}$ by the apparatus $\mathcal{A}_{2}$. However, it follows from (11) that
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. This implies that the combination of the apparatus $\mathcal{A}_{1}$ with the apparatus $\mathcal{A}_{2}$ can be regarded as a single measurement, which is described by the measurement operators $\{M^{1,2}_{m_{1},m_{2}}\}_{(m_{1},m_{2})\in\Omega_{1}\times\Omega_{2}}$ and whose all possible outcomes form the set $\Omega_{1}\times\Omega_{2}$. Thus, we can say that the measurement $\mathcal{M}^{1}$ by the apparatus $\mathcal{A}_{1}$ is really completed at the instant when the measurement $\mathcal{M}^{2}$ by the apparatus $\mathcal{A}_{2}$ is completed. The generalization of this idea to the general case where there are $n$ apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ with an arbitrary positive integer $n$ is the notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ with an arbitrary integer $i\in\{1,\dots,n\}$, introduced in Definition 7.4 above. We think that the measurement by each apparatus $\mathcal{A}_{i}$ is really completed at the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ itself.
@@
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
@@
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$.
@@
(i) Let $i\in\{1,\dots,n\}$, and let $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. Suppose that the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$. Then, in order to prove Proposition 7.5 (i), it is sufficient to show that, for every $k=i,\dots,n$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
@@
Since the state of the system $\mathcal{S}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exists a set $\{|\Upsilon_{R}[l,m_{k}]\rangle\}_{l,m_{k}}$ of vectors in $\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $l$ and every $m_{k}\in\Omega_{k}$,
@@
Since the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (16) that $M^{k}_{m_{k}}|\Gamma[l]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $l\geq 2$. Thus, using the completeness relation
@@
(ii) Let $i,j\in\{1,\dots,n\}$ with $i<j$, and let $m_{i}\in\Omega_{i}$. Suppose that the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. Then, in order to prove Proposition 7.5 (ii), it is sufficient to show that, for every $k=j,\dots,n$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
@@
Since the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exist a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $m\in\Omega_{i}$ and every $m_{k}\in\Omega_{k}$,
@@
$\begin{split}|\Psi^{k}_{\mathrm{Total}}\rangle&=(U_{k}(|\Psi\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle))\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Psi\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m\in\Omega_{i}}\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Gamma[m]\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\end{split}$ (20)
Since the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (19) that $M^{k}_{m_{k}}|\Gamma[m]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $m\in\Omega_{i}$ with $m\neq m_{i}$. Thus, using the completeness relation
@@
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m_{i}]\rangle\otimes|\Theta_{R}[m,l]\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.$

### v2

@@ 7.2 Confirming points for the final states of apparatuses
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{j}\in\Omega_{j}$, every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exists a vector $|\Theta_{R}\rangle\in\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
@@
For any $j=1,\dots,n$, we say that the states of the system $\mathcal{S}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the states of the system $\mathcal{S}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
@@
The notion of the confirming point for the states of the system $\mathcal{S}$ is defined as follows: Let $V$ be the set of all $j\in\{1,\dots,n\}$ such that the states of the system $\mathcal{S}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V\neq\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the system $\mathcal{S}$ itself if $j=1$ and the apparatus $\mathcal{A}_{j-1}$ otherwise, where $j:=\min V$. If $V=\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is defined as the apparatus $\mathcal{A}_{n}$. ∎
It is then easy to see that the notion of the confirming point for the states of the system $\mathcal{S}$ can be equivalently characterized as follows: Let $C$ be the set of all $k\in\{1,\dots,n\}$ such that the states of the system $\mathcal{S}$ are not unchanged after the measurement by the apparatus $\mathcal{A}_{k}$. If $C\neq\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is the apparatus $\mathcal{A}_{l}$ with $l:=\max C$. If $C=\emptyset$, then the confirming point for the states of the system $\mathcal{S}$ is the system $\mathcal{S}$ itself.
@@
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{j}$ if for every $m_{i}\in\Omega_{i}$, every $m_{j}\in\Omega_{j}$, every $|\Psi_{L}\rangle\in\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$, and every $|\Psi_{R}\rangle\in\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ there exist a set $\{|\Theta_{L}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Theta_{R}[l]\rangle\}_{l}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{j-1}$ such that
@@
For any $i,j=1,\dots,n$ with $i<j$, we say that the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$ if the final states of the apparatus $\mathcal{A}_{i}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{k}$ for all $k\geq j$. ∎
@@
Let $i\in\{1,\dots,n\}$. The notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as follows: Let $V_{i}$ be the set of all $j\in\{i+1,\dots,n\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. If $V_{i}\neq\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{j-1}$ with $j:=\min V_{i}$. If $V_{i}=\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is defined as the apparatus $\mathcal{A}_{n}$. Thus, the confirming point for the final states of the apparatus $\mathcal{A}_{n}$ is the apparatus $\mathcal{A}_{n}$ itself. ∎
It is then easy to see that the notion of the confirming point for the final states of an apparatus can be equivalently characterized as follows: Let $i\in\{1,\dots,n\}$, and let $C_{i}$ be the set of all $k\in\{i+1,\dots,n\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ are not unchanged after the measurement by the apparatus $\mathcal{A}_{k}$. If $C_{i}\neq\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is the apparatus $\mathcal{A}_{l}$ with $l:=\max C_{i}$. If $C_{i}=\emptyset$, then the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is the apparatus $\mathcal{A}_{i}$ itself.
Now, let us consider the case where there are only two apparatuses $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with $n=2$ and the final states of the apparatus $\mathcal{A}_{1}$ are not unchanged after the measurement by the apparatus $\mathcal{A}_{2}$. In this case, we cannot think the measurement $\mathcal{M}^{1}$ performed by the apparatus $\mathcal{A}_{1}$ to be completed immediately after the measurement $\mathcal{M}^{1}$ itself, since the state of the apparatus $\mathcal{A}_{1}$ is disturbed by the subsequent measurement $\mathcal{M}^{2}$ performed by the apparatus $\mathcal{A}_{2}$. However, it follows from (11) that
@@
for every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. This implies that the combination of the measurement $\mathcal{M}^{1}$ performed by the apparatus $\mathcal{A}_{1}$ with its subsequent measurement $\mathcal{M}^{2}$ performed by the apparatus $\mathcal{A}_{2}$ can be regarded as a single measurement, which is described by the measurement operators $\{M^{1,2}_{m_{1},m_{2}}\}_{(m_{1},m_{2})\in\Omega_{1}\times\Omega_{2}}$ and whose all possible outcomes form the set $\Omega_{1}\times\Omega_{2}$. Thus, we can say that the measurement $\mathcal{M}^{1}$ performed by the apparatus $\mathcal{A}_{1}$ is really completed at the instant when the measurement $\mathcal{M}^{2}$ performed by the apparatus $\mathcal{A}_{2}$ is completed. The generalization of this idea to the general case where there are $n$ apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ with an arbitrary positive integer $n$ is the notion of the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ with an arbitrary integer $i\in\{1,\dots,n\}$, introduced in Definition 7.4 above. We think that the measurement by each apparatus $\mathcal{A}_{i}$ is really completed at the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ itself.
@@
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
@@
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$.
@@
(i) Let $i\in\{1,\dots,n\}$, and let $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$. Suppose that the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{i}$. Then, in order to prove Proposition 7.5 (i), it is sufficient to show that, for every $k=i,\dots,n$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
@@
Since the states of the system $\mathcal{S}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exists a set $\{|\Upsilon_{R}[l,m_{k}]\rangle\}_{l,m_{k}}$ of vectors in $\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $l$ and every $m_{k}\in\Omega_{k}$,
@@
Since the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (16) that $M^{k}_{m_{k}}|\Gamma[l]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $l\geq 2$. Thus, using the completeness equation
@@
(ii) Let $i,j\in\{1,\dots,n\}$ with $i<j$, and let $m_{i}\in\Omega_{i}$. Suppose that the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$. Then, in order to prove Proposition 7.5 (ii), it is sufficient to show that, for every $k=j,\dots,n$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{k}$.
@@
Since the final states of the apparatus $\mathcal{A}_{i}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{k}$, we have that there exist a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{\mathcal{S}}\otimes\mathcal{H}_{1}\otimes\dots\otimes\mathcal{H}_{i-1}$ and a set $\{|\Upsilon_{L}[m,m_{k},h]\rangle\}_{m,m_{k},h}$ of vectors in $\mathcal{H}_{i+1}\otimes\dots\otimes\mathcal{H}_{k-1}$ such that, for every $m\in\Omega_{i}$ and every $m_{k}\in\Omega_{k}$,
@@
$\begin{split}|\Psi^{k}_{\mathrm{Total}}\rangle&=(U_{k}(|\Psi\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle))\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Psi\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle\\ &=\sum_{m\in\Omega_{i}}\sum_{m_{k}\in\Omega_{k}}(M^{k}_{m_{k}}|\Gamma[m]\rangle)\otimes|\Phi^{k}[m_{k}]\rangle\otimes|\Phi^{k+1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.\end{split}$ (20)
Since the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{k}$ by the assumption, it follows from (19) that $M^{k}_{m_{k}}|\Gamma[m]\rangle=0$ for every $m_{k}\in\Omega_{k}$ and every $m\in\Omega_{i}$ with $m\neq m_{i}$. Thus, using the completeness equation
@@
$|\Psi^{k-1}_{\mathrm{Total}}\rangle=\sum_{l}|\Theta_{L}[l]\rangle\otimes|\Phi^{i}[m_{i}]\rangle\otimes|\Theta_{R}[m_{i},l]\rangle\otimes|\Phi^{k}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle.$